A chair of weight 140 N lies atop a horizontal floor; the floor is not frictionless. You push on the chair with a force of F = 35.0 N directed at an angle of 40.0 ∘ below the horizontal and the chair slides along the floor.
Part A
Using Newton's laws, calculate n, the magnitude of the normal force that the floor exerts on the chair.
Answer #1

The forces acting downward ( ⊥ to ground) are the weight of the chair which is 140 Newton.
and the vertical component of the applied force to push the chair
= F*sin θ = 35*sin40° = 35*0.643 = 22.497 N
so, Total downward force = 140 + 22.497 = 162.497 N
As per Newton's third law
the magnitude of the normal force that the floor exerts on the chair. = 162.497 N.